OpenStudy (haifurriesftw):

You are debating about whether to buy a new car for $19,072.00 or a used car for $15,635.00. Sales tax is 4.5%. You plan to make a down payment of $1,200.00 and the APR on the loan is 7.0%. What is the difference in interest accrued by the end of the first month?

OpenStudy (haifurriesftw):

I know how to figure out all but the APR portion. 19072*1.045=19930.24 19930.24-1200=18730.24 15635*1.045=16338.575 16338.575-1200=15138.575